# How Consistent Are Humans When Grading Programming Assignments?

### v2

@@ Abstract.
We measured the inter-rater reliability of the groups using Krippendorf’s $\alpha$ – an $\alpha>0.667$ is recommended to make tentative conclusions based on the rating. Our groups were inconsistent, with an average $\alpha=0.2$ when grading correctness and an average $\alpha<0.1$ for code elegance, readability and documentation.
@@
Our results show that human graders in our study can not agree on the grade to give a piece of student work and are often individually inconsistent, suggesting that the idea of a “gold standard” of human grading might be flawed, and highlights that a shared rubric alone is not enough to ensure consistency, and other aspects such as assessor training and alternative grading practices should be explored to improve the consistency of human grading further when grading programming assignments.
@@ 2.2. Rubric Design
Jonsson and Svingby (2007) reviewed the reliability, validity and educational consequences of using rubrics in assessment. They analysed 75 studies and found that rubrics can improve the reliability of assessment if they are analytic, topic-specific, and complemented with exemplars or rater training, especially with open-ended assessment, as it is not always possible to limit the format of the assessment to achieve high levels of reliability without sacrificing the validity.
@@
To increase the agreement in teachers’ grading, Anders Jönsson and Hartell (2021) conducted a study into whether analytic or holistic rubrics enabled consistent assessment. Analytic assessment involves assessing different aspects of student performance, such as mechanics, grammar, style and organisation. On the other hand, holistic assessment means making an overall assessment that considers all criteria simultaneously (Anders Jönsson and Hartell, 2021). Anders Jönsson and Hartell’s study investigated analytic and holistic grading in both an English as a foreign language (EFL) and a mathematics high-school course. They found that analytic grading is slightly preferable to holistic assessment regarding teacher agreement for both EFL and mathematics. They further conclude that the results support the argument that teachers’ grading is ‘complex, intuitive and tacit’ (Bloxham et al., 2016) and that any attempts to achieve grading consistency are likely futile. However, their results also indicate a small step towards increasing agreement.
@@ 2.3. Teaching Assistants and Assessment
Riese and Bälter (2022) conducted a qualitative analysis into how course coordinators use assessment throughout their course and what they expect from TAs. They found that the seven experienced course coordinators in CS1 courses typically use TAs to provide continuous assessments for large cohorts. They discuss how some TAs set the bar too high while others set it too low, which can lead to consistency and that some TAs can view their students as friends, leading to ethical dilemmas. Similarly, to Kristiansen et al., Riese and Bälter suggest that not all tasks are suitable for TAs and that assessing students’ competency and knowledge can be difficult for experienced teachers.
@@ 3.3. Rubric
The rubric we provided participants can be found in Appendix B. This is the original rubric from the assignment from where the submissions were captured. The participants were asked to supply letter grades between A and F, with ‘+’ and ‘-’ grades for all letters; for exceptional submissions, participants could award A++. Participants were asked to grade the following programming skills:
@@
While code elegance, readability, and documentation are all aspects of code quality, we asked participants to provide grades for each aspect, allowing for richer data capture of different aspects of code quality. Similar aspects have been utilised in multiple different studies (Messer et al., 2024; Neutens et al., 2022)
@@ 4.1. Group Consistency (RQ1)
As many percent chance agreement-based coefficients, including Krippendorf’s $\alpha$ and Fleiss’ $\kappa$, underestimate the agreement when one of the labels is dominant (Gwet, 2008; Doughty et al., 2024; Wongpakaran et al., 2013), we also compute Gwet’s $AC_{2}$, which is a weighted variant to Gwet’s $AC_{1}$, which was designed to be robust in the high prevalence of a single class (Gwet, 2008). Gwet’s $AC_{2}$ is more robust in instances of high prevalence, as it removes from consideration all agreements that occurred by chance (Gwet, 2008).
As we planned, each assignment was graded by four of the graders rather than all graders, making Krippendorff’s $\alpha$ and Gwet’s $AC_{2}$ the ideal measures. Using Krippendorf’s $\alpha$ and Gwet’s $AC_{2}$ varies from our pre-registration, as our original choice of metric (Fliess’ $\kappa$) required a complete set of grades and did not consider the ordinal nature of grading, so we realised that this was not the best choice.
@@ 4.2. Individual Consistency (RQ2)
To evaluate the consistency of individual participants, we calculate the distance between the grade awarded in the first and second batches for each participant and for each skill. We could not use typical inter-rater analysis to evaluate self-consistency, as the two grades are too few for inter-rater reliability metrics, including Krippendorf’s $\alpha$, Gwet’s $AC_{2}$ and Fliess’ $\kappa$. We chose to limit the number of repeated assignments to ensure that the participants were unaware that they were grading a duplicate and to maximise the number of annotations for the dataset.
@@ 5.3.1. Inter-Rater Metrics
We opted to use the ordinal metric within Krippendorf’s $\alpha$, as letter grades are inherently ordinal; they are categorical and have an order. The ordinal metric applies weights to the $\alpha$ calculation by utilising the ordinal difference between ranks (Krippendorff, 2013).
When using Krippendforf’s alpha, if the raters completely agree, $\alpha=1$; if there is no agreement, then $\alpha\leq 0$. Krippendorf also suggests that $\alpha>0.800$ is acceptable and $0.800>\alpha\geq 0.667$ are only for drawing tentative conclusions (Krippendorff, 2013).
@@
Figure 6 shows the $\alpha$ for each of the seven groups and each skill. These results show that in our study, consistency was low, but were slightly more likely to agree on the grades for correctness and code elegance but did not agree on the grades for readability or documentation, with the average $\alpha$ being 0.2202, 0.0986, 0.0456, and 0.0548, respectively.
Figure 6 shows the results of Gwet’s $AC_{2}$. Following the same bounds defined by Krippendorf, no groups achieve acceptable agreement ($AC_{2}>0.800$) for any of the skills evaluated. However, some groups show partial agreement with $AC_{2}\geq 0.667$ for specific skills, though on average, no skill had an $AC_{2}\geq 0.667$, with the average $AC_{2}$ for correctness being 0.6286, for code elegance being 0.5924, for readability being 0.6383 and for documentation being 0.5966.
@@ 5.3.2. Grading Experience
Our participants had varying grading experience levels, which could have affected the consistency of their grading. Figure 7 shows the distribution of Krippendorf’s $\alpha$ for each of the four skills. Independent of the years of experience, no groups achieved an agreement close to the threshold for partial agreement. Our results for 1-2 years and 2-3 years of experience show varying levels of agreement, as the sample size of participants who had more than one year of grading experience was small, which is typical at many institutions where undergraduate, master’s and PhD students all assess student work (Mirza et al., 2019), as many students typically graduate after one or two years of working as teaching assistants. Those with one year of experience in grading show no significant improvement in inter-rate agreement compared to those with no grading experience.
Group 17-20, which had the highest average $\alpha=0.2571$ consisted of three master’s students and one third-year student. Another group with a similar composition was group 1-4, consisting of three master’s students and one PhD student; however, they had an average $\alpha=0.1225$, the fourth lowest average $\alpha$.
Group 25-28, which had the highest proportion of PhD students, at two, with one master’s and one third-year, had the second highest average $\alpha=0.1922$. Those that primarily consisted of third-year students, at two, with one master’s and one-third year, performed the worst in terms of inter-rater reliability, with Group 13-16 having an average $\alpha=-0.0324$ and Group 5-8 having an average $\alpha=-0.05981$. In most cases, being of a higher degree level when grading CS1 introductory programming does not increase the consistency when using multiple graders.
Figure 7. This figure shows the distribution of Krippendorf’s $\alpha$ for the participant’s grading experience and all four skills. The box plots’ box shows the inter-quartile range, while the whiskers extend to points within 1.5 of the inter-quartile range. The thresholds for good and partial agreement are the thresholds provided by Krippendorf (Krippendorff, 2013). The bar plot shows the proportion of participants with varying levels of grading experience.
@@ 5.4.1. Grade Range
The average absolute distance between grades when regrading the same assignment a second time was correctness = 1.7857, code elegance = 1.2500, readability = 1.3571 and documentation = 1.5714. These results show that for the consistency of the individual graders in our study, on average, they were between one and two grades apart between batches one and two, with a higher variability in correctness and documentation, and were more consistent with themselves when grading code elegance or readability.
@@ 6.1. How Consistently Does a Group of Graders Apply a Rubric? (RQ1)
Consistency was very poor overall. Even the best values achieved for Krippendorf’s $\alpha$ reflect distinctly poor consistency, with none of our seven groups of graders achieving an $\alpha>0.667$, the lower bound to consider reliable (Krippendorff, 2013), with the average $\alpha=0.10480$ across all groups and skills.
@@
However, those with one year of grading experience had similar inter-rater reliability to those with no previous grading experience, indicating that more grading experience alone will not improve the grading consistency to an acceptable level. This supports the idea that to conduct a consistent assessment, assessors require adequate training to interpret and apply a rubric (Riese and Bälter, 2022; Kristiansen et al., 2024), and that assessment is complex and intuitive (Bloxham et al., 2016; Anders Jönsson and Hartell, 2021).
Rubric design plays an important role in the validity and consistency of the rubric (Jonsson and Svingby, 2007). The rubric provided to our participants follows the standard practice for rubric design (Andrade, 2005; Reddy and Andrade, 2010) and includes evaluation criteria, the definitions of quality and a scoring strategy.
@@
While a more explicit rubric may improve the consistency of the grading, it would have reduced the validity of the assignment (Jonsson and Svingby, 2007), as a more explicit rubric would require a less open and creative assignment, limiting the students’ self-efficacy (Sharmin et al., 2019), and learning opportunities. Furthermore, even an explicit and analytic rubric may not infer consistent assessment, as grading consistently can be extremely difficult even with well-designed assessment criteria (Anders Jönsson and Hartell, 2021).
@@ 6.2. Are Individual Graders Consistent at Applying a Rubric? (RQ2)
The lower consistency of correctness could be due to many of our participants having not graded before, which is typical when using teaching assistants (Mirza et al., 2019; Riese and Bälter, 2022; Riese et al., 2021; Wald and Harland, 2018). While some may have completed the assignment as a student, they would have only seen a limited set of other submissions. Another aspect could be the participants having to learn the specificity of what makes a submission good for this specific assignment; as the assignment is open-ended many different solutions can meet the requirements to the same standard, but with the specificity of the rubric, that standard is up interpretation by the grader. Similar to multiple grader consistency, a more specific rubric could increase the consistency, but at the cost of the validity (Jonsson and Svingby, 2007) and creativity of the assignment.
@@ 7.2. External
Our raw data and participants were all captured from one institution and one course, which could limit how these results can be applied to other institutions or courses. However, as assessment is such a core aspect of education and students’ prospects, we believe our results highlight a fundamental issue with how summative assessment is currently conducted. This is further supported by many institutions implementing rigorous assessment procedures to ensure that the assessment practices within their institutions are fair to all students.
@@ 8. Conclusion
We split our 28 participants into groups of 4; each group graded the same 40 assignments in two batches of 20, with a two-week gap between them. To investigate the consistency of using multiple graders, we evaluated the inter-rater reliability using Krippendorf’s $\alpha$ (Krippendorff, 2013), which factors the ordinal nature of grades by applying weights to the metric calculation. We found that our participants provided inconsistent grades across correctness, code elegance, readability and documentation, with the most consistent being correctness with an average $\alpha=0.2$ and code elegance, readability and documentation all having an average $\alpha<0.1$ – where an $\alpha\geq 0.667$ is required to draw tentative conclusions and an $\alpha>0.8$ suggests acceptable consistency (Krippendorff, 2013). This indicates a very high level of inconsistency of grading between graders.
@@
Our study supports the notion that assessment is complex and intuitive (Bloxham et al., 2016), and that having a well-designed open-ended assignment may not imply consistent evaluation (Jonsson and Svingby, 2007), and suggests that there is a balance to be found between the openness of the assignment and the specificity of the rubric. It also highlights that further research is required to improve the consistency of a group of graders and individual graders when grading open-ended programming assignments, whether it be alternative grading practices or methods of effectively training undergraduate and graduate teaching assistants.

### v3

@@ Abstract.
We measured the inter-rater reliability of the groups using Krippendorff’s $\alpha$ – an $\alpha>0.667$ is recommended to make tentative conclusions based on the rating. Our groups were inconsistent, with an average $\alpha=0.2$ when grading correctness and an average $\alpha<0.1$ for code elegance, readability and documentation.
@@
Our results show that human graders in our study can not agree on the grade to give a piece of student work and are often individually inconsistent, suggesting that the idea of a “gold standard” of human grading might be flawed. This highlights that a shared rubric alone is not enough to ensure consistency, and other aspects such as assessor training and alternative grading practices should be explored to improve the consistency of human grading further when grading programming assignments.
@@ 2.2. Rubric Design
Jönsson and Svingby (2007) reviewed the reliability, validity and educational consequences of using rubrics in assessment. They analysed 75 studies and found that rubrics can improve the reliability of assessment if they are analytic, topic-specific, and complemented with exemplars or rater training, especially with open-ended assessment, as it is not always possible to limit the format of the assessment to achieve high levels of reliability without sacrificing the validity.
@@
To increase the agreement in teachers’ grading, Jönsson et al. (2021) conducted a study into whether analytic or holistic rubrics enabled consistent assessment. Analytic assessment involves assessing different aspects of student performance, such as mechanics, grammar, style and organisation. On the other hand, holistic assessment means making an overall assessment that considers all criteria simultaneously (Jönsson et al., 2021). Jönsson et al.’s study investigated analytic and holistic grading in both an English as a foreign language (EFL) and a mathematics high-school course. They found that analytic grading is slightly preferable to holistic assessment regarding teacher agreement for both EFL and mathematics. They further conclude that the results support the argument that teachers’ grading is ‘complex, intuitive and tacit’ (Bloxham et al., 2016) and that any attempts to achieve grading consistency are likely futile. However, their results also indicate a small step towards increasing agreement.
@@ 2.3. Teaching Assistants and Assessment
Riese and Bälter (2022) conducted a qualitative analysis into how course coordinators use assessment throughout their course and what they expect from TAs. They found that the seven experienced course coordinators in CS1 courses typically use TAs to provide continuous assessments for large cohorts. They discuss how some TAs set the bar too high while others set it too low, which can lead to inconsistency and that some TAs can view their students as friends, leading to ethical dilemmas. Similarly, to Kristiansen et al., Riese and Bälter suggest that not all tasks are suitable for TAs and that assessing students’ competency and knowledge can be difficult for experienced teachers.
@@ 3.3. Rubric
The rubric we provided participants can be found in Appendix B. This is the original rubric from the assignment from which the submissions were captured. The participants were asked to supply letter grades between A and F, with ‘+’ and ‘-’ grades for all letters; for exceptional submissions, participants could award A++. The rubric in Appendix B includes number bounds for each letter; this is due to Gradescope only supporting numerical grades. Within Gradescope, we supplied options that corresponded to the letter grades, with ‘+’ and ‘-’. The numbers were not used for this study, and participants were instructed to select letter grades in Gradescope.
Participants were asked to grade the following programming skills:
@@
While code elegance, readability, and documentation are all aspects of code quality, we asked participants to provide grades for each aspect, allowing for richer data capture of different aspects of code quality. Similar aspects have been utilised in multiple different studies (Messer et al., 2024; Neutens et al., 2022). Typically, these three aspects of code quality and correctness would be aggregated into a single final grade; we chose not to aggregate the grades, as different assignments will provide different weightings on these aspects. Thus, we analyse the consistency of grading code elegance, readability, documentation, and correctness, individually. We leave the methods of combining the grades for educators, but it does not affect this investigation into the consistency of grading the constituent components.
@@ 4.1. Group Consistency (RQ1)
As many percent chance agreement-based coefficients, including Krippendorff’s $\alpha$ and Fleiss’ $\kappa$, underestimate the agreement when one of the labels is dominant (Gwet, 2008; Doughty et al., 2024; Wongpakaran et al., 2013), we also compute Gwet’s $AC_{2}$, which is a weighted variant to Gwet’s $AC_{1}$, which was designed to be robust in the high prevalence of a single class (Gwet, 2008). Gwet’s $AC_{2}$ is more robust in instances of high prevalence, as it removes from consideration all agreements that occurred by chance (Gwet, 2008).
As we planned, each assignment was graded by four of the graders rather than all graders, making Krippendorff’s $\alpha$ and Gwet’s $AC_{2}$ the ideal measures. Using Krippendorff’s $\alpha$ and Gwet’s $AC_{2}$ varies from our pre-registration, as our original choice of metric (Fliess’ $\kappa$) required a complete set of grades and did not consider the ordinal nature of grading, so we realised that this was not the best choice.
@@ 4.2. Individual Consistency (RQ2)
To evaluate the consistency of individual participants, we calculate the distance between the grade awarded in the first and second batches for each participant and for each skill. We could not use typical inter-rater analysis to evaluate self-consistency, as the two grades are too few for inter-rater reliability metrics, including Krippendorff’s $\alpha$, Gwet’s $AC_{2}$ and Fliess’ $\kappa$. We chose to limit the number of repeated assignments to ensure that the participants were unaware that they were grading a duplicate and to maximise the number of annotations for the dataset.
@@ 5.3.1. Inter-Rater Metrics
We opted to use the ordinal metric within Krippendorff’s $\alpha$, as letter grades are inherently ordinal; they are categorical and have an order. The ordinal metric applies weights to the $\alpha$ calculation by utilising the ordinal difference between ranks (Krippendorff, 2013).
When using Krippendorff’s alpha, if the raters completely agree, $\alpha=1$; if there is no agreement, then $\alpha\leq 0$. Krippendorff also suggests that $\alpha>0.800$ is acceptable and $0.800>\alpha\geq 0.667$ are only for drawing tentative conclusions (Krippendorff, 2013).
@@
Figure 6 shows the $\alpha$ for each of the seven groups and each skill. These results show that in our study, consistency was low, but were slightly more likely to agree on the grades for correctness and code elegance but did not agree on the grades for readability or documentation, with the average $\alpha$ being 0.220, 0.099, 0.046, and 0.055, respectively.
Figure 6 shows the results of Gwet’s $AC_{2}$. Following the same bounds defined by Krippendorff, no groups achieve acceptable agreement ($AC_{2}>0.800$) for any of the skills evaluated. However, some groups show partial agreement with $AC_{2}\geq 0.667$ for specific skills, though on average, no skill had an $AC_{2}\geq 0.667$, with the average $AC_{2}$ for correctness being 0.629, for code elegance being 0.592, for readability being 0.638 and for documentation being 0.596.
@@ 5.3.2. Grading Experience
Our participants had varying grading experience levels, which could have affected the consistency of their grading. Figure 7 shows the distribution of Krippendorff’s $\alpha$ for each of the four skills. Independent of the years of experience, no groups achieved an agreement close to the threshold for partial agreement. Our results for 1-2 years and 2-3 years of experience show varying levels of agreement, as the sample size of participants who had more than one year of grading experience was small, which is typical at many institutions where undergraduate, master’s and PhD students all assess student work (Mirza et al., 2019), as many students typically graduate after one or two years of working as teaching assistants. Those with one year of experience in grading show no significant improvement in inter-rate agreement compared to those with no grading experience.
Group 17-20, which had the highest average $\alpha=0.257$ consisted of three master’s students and one third-year student. Another group with a similar composition was group 1-4, consisting of three master’s students and one PhD student; however, they had an average $\alpha=0.123$, the fourth lowest average $\alpha$.
Group 25-28, which had the highest proportion of PhD students, at two, with one master’s and one third-year, had the second highest average $\alpha=0.192$. Those that primarily consisted of third-year students, at two, with one master’s and one-third year, performed the worst in terms of inter-rater reliability, with Group 13-16 having an average $\alpha=-0.032$ and Group 5-8 having an average $\alpha=-0.0598$. In most cases, being of a higher degree level when grading CS1 introductory programming does not increase the consistency when using multiple graders.
Figure 7. This figure shows the distribution of Krippendorff’s $\alpha$ for the participant’s grading experience and all four skills. The box plots’ box shows the inter-quartile range, while the whiskers extend to points within 1.5 of the inter-quartile range. The thresholds for good and partial agreement are the thresholds provided by Krippendorff (Krippendorff, 2013). The bar plot shows the proportion of participants with varying levels of grading experience.
@@ 5.4.1. Grade Range
The average absolute distance between grades when regrading the same assignment a second time was correctness = 1.786, code elegance = 1.250, readability = 1.357 and documentation = 1.571. These results show that for the consistency of the individual graders in our study, on average, they were between one and two grades apart between batches one and two, with a higher variability in correctness and documentation, and were more consistent with themselves when grading code elegance or readability.
@@ 6.1. How Consistently Does a Group of Graders Apply a Rubric? (RQ1)
Consistency was very poor overall. Even the best values achieved for Krippendorff’s $\alpha$ reflect distinctly poor consistency, with none of our seven groups of graders achieving an $\alpha>0.667$, the lower bound to consider reliable (Krippendorff, 2013), with the average $\alpha=0.105$ across all groups and skills.
@@
However, those with one year of grading experience had similar inter-rater reliability to those with no previous grading experience, indicating that more grading experience alone will not improve the grading consistency to an acceptable level. This supports the idea that to conduct a consistent assessment, assessors require adequate training to interpret and apply a rubric (Riese and Bälter, 2022; Kristiansen et al., 2024), and that assessment is complex and intuitive (Bloxham et al., 2016; Jönsson et al., 2021).
Rubric design plays an important role in the validity and consistency of the rubric (Jönsson and Svingby, 2007). The rubric provided to our participants follows the standard practice for rubric design (Andrade, 2005; Reddy and Andrade, 2010) and includes evaluation criteria, the definitions of quality and a scoring strategy.
@@
While a more explicit rubric may improve the consistency of the grading, it could reduce the validity of the assignment (Jönsson and Svingby, 2007), as a more explicit rubric would require a less open and creative assignment, limiting the students’ self-efficacy (Sharmin et al., 2019), and learning opportunities. Furthermore, even an explicit and analytic rubric may not infer consistent assessment, as grading consistently can be extremely difficult even with well-designed assessment criteria (Jönsson et al., 2021). However, future work should investigate which aspects of code quality can be explicitly defined in the rubric and if defining these objectively improves consistency when grading open-ended assignments. Explicitly defined criteria could include explicitly defining code style conventions students should follow, such as naming conventions, avoiding single-character variable names, indentation, and spaces around operators.
@@ 6.2. Are Individual Graders Consistent at Applying a Rubric? (RQ2)
The lower consistency of correctness could be due to many of our participants having not graded before, which is typical when using teaching assistants (Mirza et al., 2019; Riese and Bälter, 2022; Riese et al., 2021; Wald and Harland, 2018). While some may have completed the assignment as a student, they would have only seen a limited set of other submissions. Another aspect could be the participants having to learn the specificity of what makes a submission good for this specific assignment; as the assignment is open-ended many different solutions can meet the requirements to the same standard, but with the specificity of the rubric, that standard is up interpretation by the grader. Similar to multiple grader consistency, a more specific rubric could increase the consistency, but at the cost of the validity (Jönsson and Svingby, 2007) and creativity of the assignment.
@@ 7.2. External
Our raw data and participants were all captured from one institution and one course, which limits how these results can be applied to other institutions or courses. Programming courses which use different rubrics, especially those that are more specific than the one provided in Appendix B, could have increased consistency. Prior work, such as by Passonneau et al. (2023) and Jönsson et al. (2021), suggests that simply making the rubric more detailed may not necessarily produce more consistent grading. Future work should evaluate how the specificity of the rubric affects consistency and the ability to provide open-ended assignments. Furthermore, different groups of graders, whether that be undergraduate, postgraduate teaching assistants, instructors, or a combination of all three, could result in varying grading consistency. Further work could investigate how different combinations of graders, training, and other methods can result in providing grades that are consistently applied between the group members.
As assessment is such a core aspect of education and students’ prospects, we believe our results highlight a fundamental issue with how summative assessment is currently conducted, whether that is caused by instructors not providing detailed rubrics or utilising multiple teaching assistants with limited training in assessment. This fundamental issue is further supported by many institutions implementing rigorous assessment procedures to ensure that the assessment practices within their institutions are fair to all students, and indicates that rubrics should be seen as potentially useful addition to the range of assessment tools available and not a solution for all concerns raised by students in terms for quality, consistency and usefulness (Cockett2018TheReview).
@@ 8. Conclusion
We split our 28 participants into groups of 4; each group graded the same 40 assignments in two batches of 20, with a two-week gap between them. To investigate the consistency of using multiple graders, we evaluated the inter-rater reliability using Krippendorff’s $\alpha$ (Krippendorff, 2013), which factors the ordinal nature of grades by applying weights to the metric calculation. We found that our participants provided inconsistent grades across correctness, code elegance, readability and documentation, with the most consistent being correctness with an average $\alpha=0.2$ and code elegance, readability and documentation all having an average $\alpha<0.1$ – where an $\alpha\geq 0.667$ is required to draw tentative conclusions and an $\alpha>0.8$ suggests acceptable consistency (Krippendorff, 2013). This indicates a very high level of inconsistency of grading between graders.
@@
While these results are only from one assignment at one institution, which utilised a generic rubric, our study supports the notion that assessment is complex and intuitive (Bloxham et al., 2016) and that having a well-designed open-ended assignment may not imply consistent evaluation (Jönsson and Svingby, 2007). Our results suggest that there is a balance to be found between the openness of the assignment and the specificity of the rubric, and future research should investigate how aspects of code quality and correctness can be explicitly defined while allowing for an open-ended assignment. It also highlights that further research is required to improve the consistency of a group of graders and individual graders when grading open-ended programming assignments, whether it be alternative grading practices or methods of effectively training undergraduate and graduate teaching assistants.